A small metal bead, labeled A, has a charge of 29 nC . It is touched to metal bead B, initially neutral, so that the two beads share the 29 nC charge, but not necessarily equally. When the two beads are then placed 5.0 cm apart, the force between them is 4.7×10−4 N . Assume that A has a greater charge.What is the charge qA and qB on the beads?
Answer #1

We know that, qA + qB = (29 x 10-9 C)

or   qA = (29 x 10-9 C) - qB { eq.1 }

According to coulomb's law, we have

F = ke qA qB / r2

(4.7 x 10-4 N) = (9 x 109 Nm2/C2) qA qB / (0.05 m)2

qA qB = (4.7 x 10-4 N) (0.05 m)2 / (9 x 109 Nm2/C2)

qA qB = (1.305 x 10-16 C2)

[(29 x 10-9 C) - qB] qB = (1.305 x 10-16 C2)

(29 x 10-9 C) qB - qB2 = (1.305 x 10-16 C2)

qB2 - (29 x 10-9 C) qB + (1.305 x 10-16 C2) = 0                                         { eq.2 }

it's an quadratic equation and we have

qB = 5.57 x 10-9 C            or          qB = 2.34 x 10-9 C

Now, using eq.1 ; qA = (29 x 10-9 C) - qB

qA = (29 x 10-9 C) - (5.57 x 10-9 C)

qA = 2.34 x 10-8 C

And

qA = (29 x 10-9 C) - (2.34 x 10-9 C)

qA = 2.66 x 10-8 C

Therefore, the charge on the beads will be given as :

qA = 2.66 x 10-8 C       and     qB = 2.34 x 10-9 C